A rectangular parcel of land is 70 ft. longer than it is wide. Each diagonal between opposite corners is 130 ft. What are the dimensions of the parcel?
Answer by nerdybill(7384)

A rectangular parcel of land is 70 ft. longer than it is wide. Each diagonal between opposite corners is 130 ft. What are the dimensions of the parcel?
.
Let w = width
then
w+70 = length
.
Applying Pythagorean theorem:
w^2 + (w+70)^2 = 130^2
w^2 + w^2+70w+70w+4900 = 16900
2w^2 + 140w + 4900 = 16900
w^2 + 70w + 2450 = 8450
w^2 + 70w - 6000 = 0
.
Factoring the left:
(w-50)(w+120) = 0
.
w = {-120, 50}
.
toss out the negative solution:
width is 50 feet
length is 50+70= 120 feet